OpenStudy (anonymous):

Solve for d: 8g+6d=41 Solve for r: C=2(3.14)r

OpenStudy (anonymous):

WILL MEDAL :-)

OpenStudy (xguardians):

I can walk you through this.

OpenStudy (xguardians):

Are the two questions separate of one another?

OpenStudy (anonymous):

They are

OpenStudy (xguardians):

Okay, I'll show you how to do the first one, then you can try the second one on your own.

OpenStudy (anonymous):

Okay thank you

OpenStudy (xguardians):

8g + 6d = 41 6d = 41 - 8g d = (41-8g)/6 Is that what you are looking for?

OpenStudy (amorfide):

if you think of 3=2+1 to solve for 2 you would subtract 1 from both sides you do this because you want to use the opposite operation so if you added 1 you would take away 1 if you multiplied you would divide and vice versa so to rearrange for 2 we do 3=2+1 3-1=2+1-1 since 1-1=0 2=2 so in your question 8g+6d=41 you want to solve for d think of 8g as one value so you want to subtract 8g from both sides to get 8g-8g+6d=41-8g so we have 6d=41-8g now we have 6d but we want d on its own since we have 6 multiplied by d we must divide both sides by 6 to get 6d/6=(41-8g)/6 since 6/6 is one we get d=(41-8g)/6 now try to do the next one